OpenStudy (anonymous):

Evaluate the expression below at x = 4. 1/x(6x up 3) a 1/24 B. 96 C. 384 D. 576

OpenStudy (anonymous):

@johnweldon1993

pooja195 (pooja195):

do u mean 6x^3

OpenStudy (johnweldon1993):

i believe she does yes

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

is it...1 / 6x^3? or just 6x^3?

pooja195 (pooja195):

6x4=24^3=72 and im sort of stuck arter that but this may help

OpenStudy (johnweldon1993):

the 3 is an EXPONENT..you dont multiply numbers and exponents :) 4^3 is equal to 4 times 4 times 4....=64

OpenStudy (anonymous):

1/x (6x up 3

OpenStudy (johnweldon1993):

hmm...so 1/4 times 6(4)^3 1/4 times 384 384 times 1 = 384 384 divided by 4 = 96 :)

OpenStudy (anonymous):

thx

OpenStudy (johnweldon1993):

anytime :)